OpenStudy (anonymous):

if a right angle triangle has a base 4 and the hypotenous =24 find the length of height=x=?

OpenStudy (bradely):

 4^2+x^2=24^2 x^2=24^2-4^2

OpenStudy (anonymous):

Answer should be 21